OpenStudy (anonymous):

What is the value for x in the proportion x over x plus ten equals seven over twelve

OpenStudy (anonymous):

\[\frac{x}{x+10}=\frac{7}{12}\iff 12x=7(x+10)\] solve for \(x\)

OpenStudy (anonymous):

14?